At December 31, 2020 and 2021, Coronado Industries had outstanding 4300 shares of $100 par value 6% cumulative preferred stock and 20000 shares of $10 par value common stock. At December 31, 2020, dividends in arrears on the preferred stock were $12100. Cash dividends declared in 2021 totaled $46000. What amounts were payable on each class of stock?
Answer #1

Preferred dividend each year=(4300*100)*6%=$25800

Hence since preferred dividend is cumulative;preferred dividend not paid for in one year is bought forward and paid for in coming years.

Hence total preferred dividend to be paid=(12100+25800)=$37900

NOTE:Any balance left over after paying preferred dividend is paid to common stockholders

Hence dividend paid to common stockholders=(46000-37900)=$8100